Find the value of x that makes the equivalent annual worth in years 1 through 10 equal to $400 per year. Use an interest rate of 10% per year.

The cash flows in year 0 and year 10 has an unknown value x, and 1 through 9 has an annual worth of $400
Answer #1

AW of cash flow = X *(A/P,10%,10) + 400 + (X-400)*(A/F,10%,10)

= X *0.162745 + 400 + (X-400)*0.062745

= X *0.162745 + 400 + X*0.062745 - 400*0.062745

= 374.902 + 0.22549 * X

As per given condition

374.902 + 0.22549 * X = 400

X = (400 - 374.902) / 0.22549 = 111.304 ~ 111.30